OpenStudy (anonymous):

Eric leans a ladder against the roof of his house so that the ladder forms a angle with the ground. The roof of the house is 13 feet above the ground. How far is the bottom of the ladder from the base of the house? Round the answer to the nearest tenth.

OpenStudy (anonymous):

are you given anything else? like an angle or another side?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is has a 70 angle

OpenStudy (anonymous):

I am not able to help you since you need more than just one variable to solve the equation \[x^2+y^2=z^2\] Where x is the distance between the bottom of the ladder and the house

OpenStudy (anonymous):

Does it say which angle is 70?

OpenStudy (anonymous):

i dont know haha thats why i need help

OpenStudy (anonymous):

Does it say which angle is 70?

OpenStudy (anonymous):

no

OpenStudy (anonymous):



OpenStudy (anonymous):

You would use tangent here which is \[\frac{ Opposite }{ Adjacent }\] \[\tan(70)=\frac{ 13 }{ x }\]

OpenStudy (anonymous):

\[x=\frac{ \tan(70) }{ 13 }\]

OpenStudy (anonymous):

.0939969168